OpenStudy (anonymous):

What is the slope of the line that passes through the points (6,2) and (4,10)? a. -1/4 b. -4 c. 4 d. 1/4

jimthompson5910 (jim_thompson5910):

Use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

(x1,y1) = (6,2) is your first point (x2,y2) = (4,10) is your second point

OpenStudy (anonymous):

\[m = \frac{ y2 - y1 }{ x2 - x1 }\] All we need to do is plug in the two points. (6,2) and (4,10) 6 is x1. 2 is y1. 4 is x2. 10 is y2. Plug them in. \[m = \frac{ 10 - 2 }{ 4 - 6 } = \frac{ 8 }{ -2 } = -4\] So, the answer is B. -4 Does this clarify things? (: